Search for Z′ in the dimuon spectrum of B → K μμ

Sascha Liechti

University of Zurich - LHCb group

Overview

1. Decay

2. Search 3. Z′ cusp 4. Outlook

Decay

• Anomalies gave rise to new search for possible lepton

flavor universality violation

• Different coupling to leptons of different generations

▷ Most to τ, least to e

▷ How can we observe different generations at once?

2